OpenStudy (anonymous):

Hector has a summer house where he resides for the months of June, July and August. Over a three-year period, he followed this schedule, except one year in which he stayed only in July and August. What fractional part of these years did Hector reside in summer house? That is my question of asking how to do a solution to the answer which is 2/9 or .22.. Please help!

OpenStudy (anonymous):

The total number of months is 3*12=36 . He stayed there in 3+3+2=8. Divide the number of months he stayed there by the total number of months: 8/36=(2*4)/(9*4)=2/9.

OpenStudy (anonymous):

Thank you again!

OpenStudy (anonymous):

It takes a machine 2minutes and 15 seconds to assembly one chair. if the machine runs continuously for 6hours how many chair will it produce?